# The Class 0 protostars in Orion: Characterizing the properties of their magnetized envelopes

### v1

@@ 4.3 B-field and Disk Formation
Observations indicate that cloud cores in star-forming environments are magnetized (e.g., Crutcher, 1999). Theoretical models suggest that B-field tension forces deflect infalling gas towards the equatorial plane, resulting in a dense, flattened pseudo-disk (e.g., Allen et al., 2003). During the collapse phase, the field lines are drawn inward into a pinched geometry (e.g., Girart et al., 2006, 2009; Stephens et al., 2013; Kwon et al., 2019), which strengthens the B-field and carries the angular momentum outward. This process is known as magnetic braking, which reduce angular momentum by one order of magnitude, potentially suppressing the formation of a centrifugally supported disk (e.g., Galli et al., 2006). However, some studies propose that an initial misalignment between the B-field and the rotation axis would lead to weaker magnetic braking of the collapsing core, enabling earlier disk formation (e.g., Joos et al., 2012; Li et al., 2013; Hirano et al., 2019; Machida et al., 2020). Recent observational studies have also reported that a disk can be formed even in young stellar objects with B-fields aligned with the rotation axis (Kwon et al., 2019; Yen et al., 2021).
@@
Huang et al. (2024) have revealed the dust polarized continuum emission and outflow signatures in protostars, finding that 40% of the protostars exhibit a mean B-field direction that is approximately perpendicular to the outflow on envelope scales. However, the relative orientation in the remaining sample appears to be random. The disk mass and disk radius taken from Tobin et al. (2020) are shown in the first row of Figure 4, which are both randomly distributed. We combine the above disk properties and envelope misalignment between the B-field and outflow, as shown in Figure 6(D–F). Protostars with complex and unresolved polarized emission, as well as those with disks significantly affected by envelope emission (HOPS-87, HOPS-317S, and HOPS-403), are excluded in our analysis. This exclusion is due to the substantial uncertainty in the averaged B-field direction or disk radius/mass for these sources. In Figure 6(D), protostars with a standard hourglass or rotated hourglass B-field structure exhibit a positive correlation between disk radius $R_{\rm disk}$ and the misalignment angle between magnetic and outflow axes $|\Delta\theta|$. In contrast, protostars with spiral field structure exhibits a random distribution. When accounting for envelopes with various B-field morphologies, a positive correlation is significant in the distribution (as shown in panel D of Figure 6). Similar discernible trends for each type of B-field morphology can be found in Figure 6(E), which illustrate a distribution between the disk mass $M_{\rm d}$ and $|\Delta\theta|$. A weaker positive correlation is noted between $M_{\rm disk}$ and $|\Delta\theta|$ when including all envelopes (standard hourglass, rotated hourglass, and spiral). This reduced correlation compared to $R_{\rm disk}$$|\Delta\theta|$ relation suggest that the derived disk mass from Tobin et al. (2020) may have potential uncertainties.
@@
Sheehan et al. (2022) found that the small disks (with radius less than $\sim 100$ au) often have their masses overestimated compared to radiative transfer model, with the discrepancy becoming more pronounced for lower-mass disks. To address this, they proposed a log-linear relation to reconcile the disk masses derived from Tobin et al. (2020) with those obtained through radiative transfer modelling, expressed as:
@@
Figure 6(F) shows the distribution between the re-estimated disk mass and $|\Delta\theta|$. The positive correlation between these variables remains apparent, suggesting that the uncertainty in disk mass derived from Tobin et al. (2020) does not significantly affect the observed correlation.
If the bottom-left point in Figure 6(D–F) is excluded, the updated correlation coefficients are $r_{p}\sim 0.54$, $r_{s}\sim 0.59$, p–value $<0.01$ for panel A, $r_{p}\sim 0.49$, $r_{s}\sim 0.53$, p–value $\sim 0.02$ for panel B, and $r_{p}\sim 0.30$, $r_{s}\sim 0.41$, p–value $\sim 0.05$, respectively. These results indicate that the correlations between disk properties and the misalignment angle remain consistent, even when the bottom-left point is excluded, suggesting that the correlation is less importantly affected by this point. The p-values, being $\lesssim$ 0.05, suggest that these correlations are statistically significant.
Huang et al. (2024) applied both the intensity-weighted method (which gives more weight to the inner protostellar envelope) and the error-weighted method (which provides an averaged weight across the selected envelope) to estimate the mean direction of the (B-field). The mean direction distributions of B-field obtained from both weighting methods were found to be comparable across the entire sample, suggesting that the measured B-field is minimally influenced by gravitational collapse and outflows. In this case, the measured B-field is likely representative of the original pre-stellar core B-field. Furthermore, the outflow direction aligns with the angular momentum of the original core. Our findings suggest that a greater misalignment between these two directions lead to weaker magnetic braking of the collapsing core and result in a more prominent disk, which is consistent with the misalignment-induced disk formation model (e.g., Joos et al., 2012).
@@ 5 Conclusions
We have calculated and analyzed the physical parameters for 55 protostelar envelopes within the BOPS sample. Non-thermal velocity dispersion is often associated with turbulence, and the polarized dust emission is thought to trace B-fields by magnetically aligned grains on envelope scales. The combination of envelope properties from the BOPS survey (Huang et al., 2024) and the disk properties from the VANDAM survey (Tobin et al., 2020) allows us to link the turbulence and B-field in the envelope to the resulting fragmentation and disk properties. The main results are as follows:
@@
    Protostellar envelopes with compact polarized dust emission tend to have a lower flux density, envelope mass, gas column density, and gas number density. While envelopes with a standard-hourglass field shape show higher values for these parameters.
@@
    Envelopes with a standard hourglass B-field structure generally show relatively low non-thermal velocity dispersion. In contrast, envelopes with a spiral B-field structure tend to have higher non-thermal velocity dispersion, indicating a more turbulent environment.
@@
    The disk mass and disk radius show no significant correlation with the different B-field structures in the envelopes. The disk inclination angles might affect the measurement of the observed B-field morphology, but the finite resolution and vertical thickness of the disks could introduce biases in the measurement of inclination angles.
@@
    The dust temperature and bolometric luminosity are not correlated with envelope masses and densities, despite these quantities being derived from luminosity. However, protostars with a significant velocity dispersion or high turbulence levels tend to exhibit increased luminosity and higher dust temperature. This correlation is likely attributed to enhanced accretion rates and energetic feedback processes, suggesting that the luminosity is predominantly driven by accretion.
@@
    The lack of correlation between fragmentation level and envelope mass is possibly due to the complex interplay of various factors influencing fragmentation beyond just the envelope mass. The fragmentation level is weakly correlated with the non-thermal velocity dispersion, suggesting that turbulence may promote envelope fragmentation on 2400 au scales. However, the non-thermal velocity dispersion may include contributions from disk rotation, envelope rotation, and/or infalling motions, necessitating further observations to confirm this scenario. Additionally, the angle dispersion of the B-field, which is inversely proportional to the field strength, is positively correlated with the fragmentation level, indicating that the pinched field suppresses the fragmentation, as it provides additional support against gravitational collapse, and/or brakes the rotation efficiently.
  6. 6.
It is important to note that our results are based on the limited spectral resolution of C17O observations, which may influence the conclusions drawn with respect to the velocity dispersion. Further observations with higher sensitivity and resolution are needed to confirm these findings and refine out conclusions.
@@
B.H. and J.M.G. acknowledge support by the grant PID2020-117710GB-I00 (MCI-AEI-FEDER, UE). B.H. also acknowledges financial support from the China Scholarship Council (CSC) under grant No. 202006660008. This work is also partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. M.F.L. acknowledges support from the European Research Executive Agency HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Skłodowska-Curie grant agreement number 101086388 (LACEGAL). M.F.L. also acknowledges the warmth and hospitality of the ICE-UB group of star formation. L.W.L. acknowledges support by NSF AST-1910364 and NSF AST-2307844. Z.Y.L. is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199. W.K. is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00342488). The authors acknowledge Anaëlle Maury for helpful discussions, Jacob Labonte for early analysis of the BOPS data. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2019.1.00086. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.

### v2

@@ 4.3 B-field and Disk Formation
Observations indicate that cloud cores in star-forming environments are magnetized (e.g., Crutcher, 1999). Theoretical models suggest that B-field tension forces deflect infalling gas towards the equatorial plane, resulting in a dense, flattened pseudo-disk (e.g., Allen et al., 2003). During the collapse phase, the field lines are drawn inward into a pinched geometry (e.g., Girart et al., 2006, 2009; Stephens et al., 2013; Kwon et al., 2019), which strengthens the B-field and carries the angular momentum outward. This process is known as magnetic braking, which reduce angular momentum by one order of magnitude, potentially suppressing the formation of a centrifugally supported disk (e.g., Galli et al., 2006). Figure 6(D) presents the distribution between the disk size $R_{\rm disk}$ and the B-field angle dispersion ($\delta\phi$, which is inversely proportional to the B-field strength) for 11 protostars with good detection in the disk radius. We find that protostellar envelopes with a standard hourglass B-field structure tend to have a smaller disk radius and angle dispersion of the B-field than those with a rotated hourglass field structure, suggesting that magnetic braking is more effective for the standard hourglass field shape. Some studies propose that an initial misalignment between the B-field and the rotation axis would lead to weaker magnetic braking of the collapsing core, enabling earlier disk formation (e.g., Joos et al., 2012; Li et al., 2013; Hirano et al., 2019; Machida et al., 2020).
@@
Huang et al. (2024) have revealed the dust polarized continuum emission and outflow signatures in protostars, finding that 40% of the protostars exhibit a mean B-field direction that is approximately perpendicular to the outflow on envelope scales. However, the relative orientation in the remaining sample appears to be random. We combine the disk radius taken from Tobin et al. (2020) and envelope misalignment between the B-field and outflow for BOPS protostars, as shown in Figure 6(E). Protostars with complex and unresolved polarized emission, as well as those with disks significantly affected by envelope emission and other component of disk emission (see Table LABEL:Tab:parameters in Appendix A and figures in Appendix B) are excluded in our analysis. This exclusion is due to the substantial uncertainty in the averaged B-field direction or disk radius/mass for these sources. Then the number of sources used in Figure 6(E) and 6(F) is 16. In panel E, protostars with standard hourglass, rotated hourglass, and spiral B-field structures exhibit a strong positive correlation between disk radius $R_{\rm disk}$ and the misalignment angle between magnetic and outflow axes $|\Delta\theta|$.
@@
Sheehan et al. (2022) found that the disk masses derived from Tobin et al. (2020) are overestimated compared to radiative transfer model, with the discrepancy becoming more pronounced for lower-mass disks. To address this, they proposed a log-linear relation to reconcile the disk masses derived from Tobin et al. (2020) with those obtained through radiative transfer modelling, expressed as:
@@
Figure 6(F) shows the distribution between the re-estimated disk mass and $|\Delta\theta|$. A similar discernible trend can be found from this panel that $M_{\rm disk}$ is positively correlated with $|\Delta\theta|$, but the correlation is weaker compared to the $R_{\rm disk}$-$|\Delta\theta|$ relation. If the bottom-left point in Figure 6(E–F) is excluded, the updated correlation coefficients are $r_{p}\sim 0.74$, $r_{s}\sim 0.75$, p–value $<0.01$ for panel E, and $r_{p}\sim 0.62$, $r_{s}\sim 0.59$, p–value $\sim 0.02$ for panel F, respectively. These results indicate that the correlations between the disk radius/mass and the misalignment angle remain consistent, even when the bottom-left point is excluded, suggesting that the correlation is less important affected by this point. The p-values, which are $<$ 0.05, suggest that these correlations are statistically significant. In addition, it is important to note from the distribution of both panel E and panel F that protostars exhibiting standard hourglass B field tend to have lower disk size, compared to those exhibiting spiral and rotated hourglass field structure.
Huang et al. (2024) applied both the intensity-weighted method (which gives more weight to the inner protostellar envelope) and the error-weighted method (which provides an averaged weight across the selected envelope) to estimate the mean direction of the (B-field). The mean direction distributions of B-field obtained from both weighting methods were found to be comparable throughout the sample, suggesting that the measured B-field is minimally influenced by gravitational collapse and outflows. In this case, the measured B-field is likely representative of the original pre-stellar core B-field. Furthermore, the outflow direction aligns with the angular momentum of the original core. Our findings suggest that a greater misalignment between these two directions lead to weaker magnetic braking of the collapsing core and result in a more prominent disk, which is consistent with the misalignment-induced disk formation model (e.g., Joos et al., 2012).
@@ 5 Conclusions
We have calculated and analyzed the physical parameters for 61 protostelar envelopes within the BOPS sample. The combination of the envelope properties derived from the BOPS (Huang et al., 2024) and the disk properties derived from the VANDAM survey (Tobin et al., 2020) allows us to link the turbulence and B-field in the envelope with the resulting fragmentation and disk properties. The main results are as follows:
@@
    Protostellar envelopes exhibiting standard-hourglass field shape show higher values of envelope mass, gas column density, and gas number density than those with other field configurations, lower non-thermal velocity dispersion have also be observed in this type. In contrast, envelopes with a spiral B-field structure tend to have higher non-thermal velocity dispersion, indicating a more turbulent environment.
@@
    The angle dispersion of the B-field, which is inversely proportional to the B-field strength, is positively correlated with the fragmentation level, suggesting that the pinched field suppresses the fragmentation, as it provides additional support against gravitational collapse, and/or brakes the rotation efficiently. However, the lack of correlations between the fragmentation level and the envelope mass, and between the fragmentation level and the non-thermal velocity dispersion, may be due to the complex interplay of various factors influencing fragmentation beyond just the envelope mass, or to the limited spectral resolution.
@@
    Protostars with standard hourglass B-field structure have smaller disk size and angle dispersion of the magnetic field than those with other field configurations, specially rotated hourglasses, but also spirals and others. This suggests that in the case of the standard hourglass field structure, the magnetic field strength is stronger, ultimately leading to a smaller disk through the more effective magnetic braking, in line with theory predictions.
@@
    The dust temperature and bolometric luminosity are not correlated with envelope masses and densities, probably because the cloud with low bolometric luminosity provides shielding from the interstellar radiation field. In addition, no clear correlation can be found between non-thermal velocity dispersion and luminosity on envelope scales, contrast with the simulation models predicted. Deeper studies on individual cores or higher-resolution observations might reveal more localized correlations that are otherwise masked by the larger-scale dynamics.
Insignificant correlations found for some pairs of parameters, especially in the analysis of the non-thermal velocity dispersion, may be due to the limited sensitivity and resolution of the observations. It is important to note that the non-thermal velocity dispersion may include contributions from disk rotation, envelope rotation, and/or infalling motions, further observations with higher sensitivity and resolution should be carried out to distinguish them and further refine our conclusions about the non-thermal velocity dispersion or the turbulence.
@@
B.H. and J.M.G. acknowledge support by the grant PID2020-117710GB-I00 (MCI-AEI-FEDER, UE). B.H. also acknowledges financial support from the China Scholarship Council (CSC) under grant No. 202006660008. This work is also partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. M.F.L. acknowledges support from the European Research Executive Agency HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Skłodowska-Curie grant agreement number 101086388 (LACEGAL). M.F.L. also acknowledges the warmth and hospitality of the ICE-UB group of star formation. L.W.L. acknowledges support by NSF AST-1910364 and NSF AST-2307844. Z.Y.L. is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199. W.K. is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00342488). The authors acknowledge Anaëlle Maury and Chat Hull for helpful discussions, Jacob Labonte for early analysis of the BOPS data. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2019.1.00086. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.
Appendix A Table
Table 2: Some parameters for the BOPS sample
Name R.A. Decl. $D$ $M_{\rm disk}$ $R_{\rm disk}$ $\langle\Theta_{\rm max}\times\Theta_{\rm min}\rangle_{\rm Decon.}$ $\theta_{\rm Inc}$ $\Delta\theta$
(h:m:s) (d:m:s) (pc) (M) (au) (${}^{\prime\prime}\times^{\prime\prime}$) () ()
Standard
HOPS-11 05:35:13:43 -05:57:57.91 388.3 135.3 $\pm$ 2.4 41.9 $\pm$ 10.0 0.13 $\times$ 0.12 22.6 14.0 $\pm$ 9.7
HOPS-87Naafootnotemark: 05:35:23.42 -05:01:30.57 392.7 / / / / /
HOPS-359 05:47:24.84  00:20:59.39 429.4 89.8 $\pm$ 2.0 62.0 $\pm$ 10.0 0.17 $\times$ 0.15 28.1 15.6 $\pm$ 7.2
HOPS-395 05:39:17.09 -07:24:24.59 397.2 16.9 $\pm$ 0.9 7.1 $\pm$ 10.0 0.02 $\times$ 0.02 0 0.1 $\pm$ 9.0
HOPS-400 05:42:45.26 -01:16:13.89 415.4 106.6 $\pm$ 4.7 48.7 $\pm$ 10.0 0.14 $\times$ 0.13 21.3 14.3 $\pm$ 6.6
HOPS-407 05:46:28.25  00:19:27.95 419.1 83.2 $\pm$ 2.4 64.1 $\pm$ 10.0 0.18 $\times$ 0.14 38.9 16.5 $\pm$ 7.8
OMC1N-8-Nccfootnotemark: 05:35:18.20 -05:20:48.57 392.8 / / / / /
Rotated
HH270IRS 05:51:22.72  02:56:05:00 405.7 161.3 $\pm$ 3.4 92.3 $\pm$ 10.0 0.27 $\times$ 0.14 58.8 68.3 $\pm$ 9.6
HOPS-78 05:35:25.96 -05:05:43.38 392.8 187.9 $\pm$ 2.1 147.1 $\pm$ 10.0 0.44 $\times$ 0.14 71.4 86.0 $\pm$ 6.6
HOPS-168bbfootnotemark: 05:36:18.94 -06:45:23.58 383.3 / / / / /
HOPS-169 05:36:36.16 -06:38:54.41 384.0 118.5 $\pm$ 2.5 53.8 $\pm$ 10.0 0.17 $\times$ 0.13 40.1 51.3 $\pm$ 7.2
HOPS-288bbfootnotemark: 05:39:56.00 -07:30:27.62 405.5 / / / / /
HOPS-310 05:42:27.68 -01:20:01.35 414.3 263.8 $\pm$ 5.3 101.3 $\pm$ 10.0 0.29 $\times$ 0.15 58.9 89.9 $\pm$ 6.4
HOPS-317Saafootnotemark: 05:46:08.38 -00:10:43.59 427.1 / / / / /
HOPS-370 05:35:27.64 -05:09:34.40 392.8 114.6 $\pm$ 2.2 112.8 $\pm$ 10.0 0.34 $\times$ 0.11 71.1 86.8 $\pm$ 5.9
HOPS-401 05:46:07.73 -00:12:21.31 426.9 185.9 $\pm$ 4.2 74.7 $\pm$ 10.0 0.21 $\times$ 0.16 40.4 78.8 $\pm$ 7.6
HOPS-409 05:35:21.37 -05:13:17.88 392.8 110.7 $\pm$ 1.1 112.1 $\pm$ 10.0 0.34 $\times$ 0.12 69.3 73.3 $\pm$ 6.1
OMC1N-4-5-ESccfootnotemark: 05:35:15.97 -05:20:14.23 392.8 / / / / /
Spiral
HOPS-96 05:35:29.72 -04:58:48.63 392.7 135.0 $\pm$ 2.3 41.4 $\pm$ 10.0 0.12 $\times$ 0.11 23.6 52.4$\pm$ 6.2
HOPS-124ababfootnotemark: 05:39:19.91 -07:26:11.22 398.0 / / / / /
HOPS-182 05:36:18.79 -06:22:10.24 385.1 35.7 $\pm$ 1.3 36.3 $\pm$ 10.0 0.11 $\times$ 0.06 56.9 32.7 $\pm$ 6.3
HOPS-303 05:42:02.65 -02:07:45.94 410.0 206.8 $\pm$ 4.4 82.9 $\pm$ 10.0 0.24 $\times$ 0.19 37.7 53.9 $\pm$ 12.2
HOPS-361Nbbfootnotemark: 05:47:04.63  00:21:47.82 430.4 / / / / /
HOPS-361S 05:47:04.78  00:21:42.79 430.4 138.6 $\pm$ 1.4 96.1 $\pm$ 10.0 0.26 $\times$ 0.15 54.8 40.2 $\pm$ 11.0
HOPS-384bbfootnotemark: 05:41:44.14 -01:54:46.00 409.5 / / / / /
HOPS-403aafootnotemark: 05:46:27.91 -00:00:52.15 428.2 / / / / /
HOPS-404aafootnotemark: 05:48:07.72  00:33:51.81 430.1 / / / / /
Others
HOPS-12W 05:35:08.63 -05:55:54.65 388.6 20.1 $\pm$ 0.4 20.5 $\pm$ 10.0 0.06 $\times$ 0.05 33.6 /
HOPS-87Sccfootnotemark: 05:35:23.67 -05:01:40.27 392.7 / / / / /
HOPS-88 05:35:22.47 -05:01:14.33 392.7 70.5 $\pm$ 1.2 34.0 $\pm$ 10.0 0.10 $\times$ 0.09 25.8 /
HOPS-164 05:37:00.43 -06:37:10.91 385.0 133.5 $\pm$ 2.8 70.6 $\pm$ 10.0 0.22 $\times$ 0.14 50.5 /
HOPS-224 05:41:32.07 -08:40:09.82 440.3 236.6 $\pm$ 5.0 107.7 $\pm$ 10.0 0.29 $\times$ 0.18 51.6 /
HOPS-250 05:40:48.85 -08:06:57.11 428.5 133.9 $\pm$ 2.9 117.2 $\pm$ 10.0 0.32 $\times$ 0.19 53.6 /
HOPS-317N 05:46:08.60 -00:10:38.49 427.1 51.6 $\pm$ 1.8 34.5 $\pm$ 10.0 0.10 $\times$ 0.09 25.8 /
HOPS-325 05:46:39.20  00:01:12.30 428.5 58.6 $\pm$ 1.1 34.9 $\pm$ 10.0 0.10 $\times$ 0.09 25.8 /
HOPS-341 05:47:00.92  00:26:21.50 430.9 28.3 $\pm$ 1.1 54.5 $\pm$ 10.0 0.15 $\times$ 0.09 53.1 /
HOPS-373E 05:46:31.10 -00:02:33.05 428.1 79.2 $\pm$ 1.8 69.8 $\pm$ 10.0 0.19 $\times$ 0.17 26.5 /
HOPS-373W 05:46:30.91 -00:02:35.16 428.1 70.3 $\pm$ 0.8 36.0 $\pm$ 10.0 0.10 $\times$ 0.09 25.8 /
HOPS-398 05:41:29.41 -02:21:16.39 408.0 271.8 $\pm$ 6.0 76.9 $\pm$ 10.0 0.22 $\times$ 0.19 30.3 /
HOPS-399 05:41:24.93 -02:18:06.66 407.9 196.7 $\pm$ 4.3 66.2 $\pm$ 10.0 0.19 $\times$ 0.15 37.9 /
HOPS-402aafootnotemark: 05:46:10.03 -00:12:16.92 426.9 / / / / /
HOPS-408 05:39:30.90 -07:23:59.75 398.9 124.5 $\pm$ 3.6 93.5 $\pm$ 10.0 0.28 $\times$ 0.24 31.0 /
OMC1N-4-5-ENccfootnotemark: 05:35:16.05 -05:20:05.75 392.8 / / / / /
OMC1N-6-7ccfootnotemark: 05:35:15.70 -05:20:39.31 392.8 / / / / /
OMC1N-8-Sccfootnotemark: 05:35:18.00 -05:20:55.77 392.8 / / / / /
Scattering
HH212M 05:43:51.41 -01:02:53.20 413.2 102.9 $\pm$ 2.3 78.6 $\pm$ 10.0 0.22 $\times$ 0.10 63.0 /
HOPS-10 05:35:09.05 -05:58:26.89 388.2 52.5 $\pm$ 1.0 46.8 $\pm$ 10.0 0.14 $\times$ 0.07 60.0 /
HOPS-12Ebbfootnotemark: 05:35:08.95 -05:55:54.99 388.6 / / / / /
HOPS-50 05:34:40.92 -05:31:44.74 391.5 172.8 $\pm$ 3.3 82.8 $\pm$ 10.0 0.25 $\times$ 0.13 58.7 /
HOPS-53 05:33:57.40 -05:23:30.00 390.5 33.4 $\pm$ 0.5 21.6 $\pm$ 10.0 0.07 $\times$ 0.05 44.4 /
HOPS-60 05:35:23.29 -05:12:03.45 392.8 152.7 $\pm$ 2.7 62.7 $\pm$ 10.0 0.19 $\times$ 0.11 54.6 /
HOPS-81 05:35:28.01 -05:04:57.36 392.8 66.4 $\pm$ 1.2 32.4 $\pm$ 10.0 0.10 $\times$ 0.06 53.1 /
HOPS-84 05:35:26.56 -05:03:55.15 392.8 159.2 $\pm$ 3.0 87.4 $\pm$ 10.0 0.26 $\times$ 0.12 62.5 /
HOPS-153 05:37:57.03 -07:06:56.27 387.9 103.4 $\pm$ 2.1 129.1 $\pm$ 10.0 0.39 $\times$ 0.10 75.1 /
HOPS-203Nbbfootnotemark: 05:36:22.87 -06:46:06.63 383.5 / / / / /
HOPS-203S 05:36:22.90 -06:46:09.54 383.5 33.1 $\pm$ 0.4 30.9 $\pm$ 10.0 0.10 $\times$ 0.06 53.1 /
HOPS-247 05:41:26.19 -07:56:51.90 430.9 227.0 $\pm$ 4.6 85.3 $\pm$ 10.0 0.23 $\times$ 0.16 45.9 /
HOPS-340 05:47:01.31  00:26:23.05 430.9 59.9 $\pm$ 1.0 41.7 $\pm$ 10.0 0.11 $\times$ 0.07 50.5 /
HOPS-354 05:54:24.26  01:44:19.87 355.4 30.3 $\pm$ 0.6 24.4 $\pm$ 10.0 0.08 $\times$ 0.07 29.0 /
HOPS-358 05:46:07.26 -00:13:30.25 426.8 110.7 $\pm$ 1.0 136.3 $\pm$ 10.0 0.38 $\times$ 0.10 74.7 /
HOPS-383 05:35:29.79 -04:59:50.38 392.8 96.0 $\pm$ 1.9 61.7 $\pm$ 10.0 0.18 $\times$ 0.12 48.2 /
Column 2 to 9 present the R.A., Decl., distance ($D$), disk mass ($M_{\rm disk}$), disk radius ($R_{\rm disk}$), deconvolved disk size ($\langle r_{\rm max}\rangle\times\langle r_{\rm min}\rangle$), disk inclined angle ($\theta_{\rm Inc}=\arccos(\langle\Theta_{\rm min}\rangle/\langle\Theta_{\rm max}\rangle)$), and the difference between outflow and B-field direction ($\Delta\theta$), respectively. The source distance, disk mass, disk radius, disk size and their uncertainty are obtained from Tobin et al. (2020). The difference between outflow and B-field is obtained from Huang et al. (2024), which is the alignment between the mean outflow direction and the uncertainty-weighted mean B-field direction.aafootnotemark: The disk emission is contaminated by the envelope dust emission.bbfootnotemark: The disk emission is contaminated by other disk emission.ccfootnotemark: No information on disk properties presented.
Appendix B Figures
B.1 Self-scattering dominated protostars
Left panels: the velocity field in color scale (obtained from the C17O (3–2) line) overlaid with the 870 $\mu$m dust polarization segments and Stokes I contours. Blue/red contours indicate the blue-/red-shifted outflow, with counter levels set at 5 times the outflow rms × (1, 2, 4, 8, 16, 32). The black contour levels for the Stokes I image are 10 times the rms × (1, 2, 4, 8, 16, 32, 64, 128, 256, 512). For the velocity field, regions with an S/N less than 4 have been flagged. The purple segments represent the polarization. The magenta dashed circle corresponds to scales with a radius of 1200 au. Right panels: 9 mm dust maps obtained from Tobin et al. (2020).
B.2 Aligned dust grain dominated protostars
Left panels: the velocity field in color scale (obtained from the C17O (3–2) line) overlaid with the 870 $\mu$m dust polarization segments and Stokes I contours. Blue/red contours indicate the blue-/red-shifted outflow, with counter levels set at 5 times the outflow rms × (1, 2, 4, 8, 16, 32). The black contour levels for the Stokes I image are 10 times the rms × (1, 2, 4, 8, 16, 32, 64, 128, 256, 512). For the velocity field, regions with an S/N less than 4 have been flagged. The white segments represent the magnetic fields. The dashed magenta circle corresponds to scales with a radius of 1200 au. The purple curves in categories of standard hourglass, rotated hourglass, and spiral indicate the B-field patterns. Right panels: 9 mm dust maps obtained from the VANDAM survey (Tobin et al., 2020).
Standard hourglass [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
Rotated hourglass [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]
Spiral [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
Others
[Uncaptioned image]              [Uncaptioned image]              [Uncaptioned image]
Others (continue) [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]